# Premerger phenomena in neutron-star binary coalescences

### v2

@@ II.1.3 SLy family
The SLy EOS is based on a semi-empirical approach that parameterises the interactions between nucleons (protons and neutrons) in terms of density-dependent coefficients of the nuclear interactions [79, 80, 127, 173]. It is derived from the Lyon-Skyrme energy-density functional (SLy) and uses a parameter set of coefficients that are chosen to accurately describe the properties of both symmetric nuclear matter (equal numbers of protons and neutrons) and neutron-rich matter. This parameter set (e.g. with terms weighting the kinetic energy of nucleons) has been fitted to various properties of finite nuclei and nuclear matter, including binding energies, charge radii, and neutron-skin thicknesses. Neutron star models constructed using the SLy EOS can reach a maximum mass of $\sim 2.1M_{\odot}$ and have a radius of $11\lesssim R\lesssim 13$ km for typical neutron stars with a mass of 1.4 $M_{\odot}$, it is therefore a relatively stiff EOS. The SLy EOS is designed to be valid for a wide range of densities, ranging from the crust up to a few times the nuclear saturation density. It agrees well with observational data and is widely used in simulations of supernovae and neutron star mergers, especially updated versions with finite-temperature corrections [173, 349, 326].
@@ II.1.4 BSk family
The BSk EOS, consisting of a series of models [308, 306, 318, 166, 307], is developed through the use of the Brussels-Skyrme (BSk) energy-density functionals [^3] The table of BSk EOS can be generated by the resources provided in http://www.ioffe.ru/astro/neutronstarG/BSk/.. Like SLy, BSk is based on Skyrme-type energy-density functionals that use density-dependent coefficients to parameterise nucleon-nucleon interactions. These functionals are fitted to match experimental data on finite nuclei and nuclear matter, resulting in various BSk models (BSk19–26, with labelling similar to that of Argonne described earlier with the appended number representing how many terms were utilised via a Hartree-Fock-Bogoliubov mass model) with parameters chosen to provide a good fit to empirical data. Like other EOS, it also assumes beta equilibrium and maintains charge neutrality, though finite-temperature corrections are included. The predicted mass-radius relationship varies depending on the BSk model chosen, but typically predict a radius of 11–13 km for 1.4 $M_{\odot}$ neutron star models, as for SLy. The BSk models also predict a relatively high maximum mass of between 2 and 2.4 $M_{\odot}$. In Fig. 2 we show the $M$-$R$ relation for one member of this family, BSk25, for which the maximum mass is 2.23 $M_{\odot}$ and the radius of the model with $1.4\,M_{\odot}$ is $\approx$ 12.3 km.
@@ II.2 Rotation and binary alignment
Angular momentum is ubiquitous in nature. In the case of neutron stars, approximate angular momentum conservation during the supernova process implies that proto-stars can be born spinning fairly quickly, even for relatively slow progenitors. Depending on the progenitor and the impact of fallback accretion [368], some neutron stars may be born with millisecond spin periods [140]; though see also Varma et al. 414, Varma and Müller 413.
@@
Even for “slow” objects (e.g. magnetars performing a full revolution at rates less than once per few seconds), it is essential to account for rotation when considering electromagnetic observables. Depending on the rotational phase the signal can be broadened or lost due to beaming effects, which impacts interpretation. Observationally, neutron stars seem to be anything between practically static up to spins of $\sim 800$ Hz. Why there is such an observational upper limit is a topic of active research, invoking various explanations from GW-enhanced spindown to centrifugal propellering either from a companion or fallback [see, e.g. 304, for a discussion].
@@
One would like to know what the likelihood of having a “rapidly” rotating star taking part in a merger is. A review of spin properties in binaries, and the formation of double neutron-star systems in general, can be found in Tauris et al. 395. The situation is somewhat complicated however, and it has been argued that there may be three sub-populations of double neutron-star systems with distinct spin distributions, pertaining to (i) short-period and low-eccentricity binaries, (ii) wide binaries, and lastly (iii) short-period, high-eccentricity binaries [36]. The canonical formation channel involves a symbiotic binary (i.e. without dynamical captures) where there are two supernovae. These supernovae are separated by a timescale that is sensitive to a number of evolutionary specifics related to common-envelope evolution and how susceptible the companion is to Roche-lobe overflow, which in turn depends primarily on the mass ratio [e.g. 412].
@@
Prior to double neutron-star formation, accretion by the first born from the non-degenerate star can be either of a disk or wind-fed nature, and will tend to spin-up the neutron star thereby leading to a “recycled” object with a spin that is (close to) aligned with the orbital angular momentum. A discussion on predictions for alignment angles can be found in Section 2.3 of Kuan et al. 232 and references therein: several binaries have non-negligible misalignment angles as measured through geodetic precession and optical polarimetric measurements [e.g. PSR B1534+12 in a double neutron-star system with a misalignment of $\sim 27^{\circ}$; 144]. Such misalignment is relevant for the excitation of non-axisymmetric modes during inspiral (see Sec. IV.3). Either way, for a spin-up rate of $\dot{\nu}\gtrsim 10^{-14}\text{ Hz s}^{-1}$, not unusual from observations or torque theory [162], the first-born neutron star could attain spins of $\gtrsim 100$ Hz within significantly sub-Hubble times $\lesssim 10^{2}$ Myr [see table 1 in 395]. Given that the dipolar magnetic field may be buried by large factors via epochs of mass infall [267, 418, 388, 389, see also Sec. II.3], the spindown that sets in, once the secondary collapses and stops gifting material, may not be sufficient to erase spinup before merger.
@@
Zhu et al. 449 estimate that between $15$ and $30\%$ binary neutron-stars will have spins measurable via GWs at $\gtrsim 90\%$ confidence. This implies the importance of accounting for spin in tidal modelling, as all mode eigenfrequencies are skewed by even a modest degree of rotation; see equation (24) and also Appendix A of Suvorov et al. 386.
@@ II.3 Magnetic fields
Although a complete survey of how the magnetic field impresses on observables associated with the neutron star population at large lies well beyond the scope of this Review [see 132, for a thorough exposition], many observations have proven that magnetic multipoles and topologically complicated structures are a reality:
@@
    The morphology of pulsed emissions are highly varied, with some systems displaying long term epochs of nulling or interpulses. Interpulse phenomenology in radio pulsars can be qualitatively explained by an oblique rotator with a multipolar magnetic field, as the emissions are then composed of multiple components [46]. The X-ray light curve from the magnetar SGR1900+14 also displays interpulse-like phenomena, which may be contributed by multiple hotspots on the surface [53]. Zhang et al. 447 suggest that starspots (i.e. localised, multipolar fields) may emerge through Hall evolution near the poles of neutron stars that are hovering around the death line, sporadically allowing the hosts to pulse and possibly explaining nulling [see also 387].
@@
    Current bundles injected into the magnetosphere by crustal motions twist the fields there, inducing multipolarity [50]. Mdels of neutron star activity often invoke crustal failures as seeding events for outbursts [such as glitches or flares, e.g. 48, 381, 103], and therefore such injections may be common. In fact, crust failures are one of the mechanisms proposed for premerger precursors; see Sec. VI.
Aside from poloidal multipoles, the neutron star magnetic field also likely contains a toroidal component:
@@
    Precession in magnetars, such as 4U 0142+61 [255] and SGR 1900+14 [256], are most straightforwardly explained through a (sub-)crustal toroidal field of strength $\gtrsim 10^{16}$ G: such a field introduces a prolate distortion along the magnetic axis, which then becomes misaligned with the rotation axis, causing free precession [see also 260, 382].
@@
    Mixed fields are necessary for the stability of the star. Magnetohydrodynamic (MHD) models and stability analyses demonstrate that purely poloidal/toroidal fields are unstable over even short timescales [397, 49, 14, 182].
@@
    A universal feature of magnetothermal evolutions is that energy is regularly swapped between the toroidal and poloidal sectors [e.g. 316, 111]. Similarly, high-order multipoles are inevitable consequences of cascade phenomena via Hall drift [387] or plastic flows [168].
@@
Figure 3 illustrates the conventional picture of pulsar operation and the presence of magnetic substructure generally (including the likelihood of internal multipoles and a superconductor; see Sec. III.4). Interpretation of neutron-star phenomena is complicated by the fact that emissions must traverse the interstellar (or intergalactic) medium, which can lens photons or GWs [see, e.g. 376].
@@
If we ignore plastic flows [though see, e.g. 169, 168, 391], the overall crustal field strength evolution may be written, after solving a simplified volume-averaged version of the induction equation following Aguilera et al. 12, as
@@
with $\tau_{\rm H}=4\pi en_{e}L^{2}/cB_{0}$ and $\tau_{\Omega}=4\pi\sigma_{\rm e}L^{2}/c^{2}$ for length-scale $L\lesssim R$ (smaller for higher-order multipoles), initial field strength $B(0)=B_{0}$, electron number-density $n_{e}$, and electrical conductivity $\sigma_{\rm e}$. For typical parameter choices in the crust, we may have therefore that $\tau_{H}\sim\text{few}\times 10^{4}(B_{0}/10^{14}\text{ G})^{-1}$ yr and $\tau_{\Omega}\sim\text{few}\times$ Myr. Given, however, that the GW inspiral time is orders of magnitude longer than either of these times unless the field strength is $\sim 10^{10}$ G, one anticipates weak fields at time of merger. This has important implications for the precursor flares we discuss in Sec. V.
It is important to note however that there are some ways around the decay issue, such as via Hall-stalling [i.e. the system reaches a quasi-equilibrium where $\tau_{\rm H}$ tends to some much larger number; 167], an absence of crustal impurities [which increase the conductivity and hence the Ohmic time; 196], plastic flows [which can work to halt electron flows and thus suppress cascading and the formation of multipoles; 168, 391], or dynamical capture (thereby avoiding issues related to decay altogether). Whether it is at all possible for magnetar-level fields to persist over cosmological timescales remains an active area of research [see, e.g. 196, 378, for discussions].
@@ III.1 Basic elements of crustal physics
Although a neutron star may be born with temperatures in excess of $\gtrsim$ MeV, especially in a merger, the system begins to cool rapidly through a sequence of beta decays (Urca reactions) as neutrinos flood outward from what is currently an envelope [e.g. 131, 339, 351]. The details of these very early stages in the star’s life are a matter of active research and lie beyond the scope of this Review; see, for instance, Sarin and Lasky 345. Still, this early behaviour is important as concerns some early post-merger phenomena within our purview, so some details are touched on in subsequent sections.
@@
where $Z$ denotes the number of protons in the ion, $a_{i}$ is the ion sphere radius (i.e. the Wigner-Seitz cell radius) such that $34\pi a_{i}^{3}/$ 3 equals the volume per ion, $1/n_{i}$, $e$ is the elementary charge, $k_{B}$ is the Boltzmann constant, and finally $T$ represents the temperature. Aside from fundamental constants, each ingredient defining $\Gamma$ varies with depth in a complicated way [82]. Once the temperature drops sufficiently ($\sim 10^{8}$ K) such that ${\Gamma}\gtrsim 175$ [e.g. 137], the liquid envelope solidifies via a first-order phase transition into an elastic material — the crust.
In general, because of the density dependence in expression (2), the crust may not entirely encompass the final $\gtrsim 1$ km of the star. It is expected at least that there will be an ‘ocean’ separating the crust from the magnetosphere [e.g. 159, see also Sec. VI.4.3], the physics of which depends strongly on temperature, meaning first of all that a zero-temperature EOS cannot apply [171] but also that no stress can be supported there. Roughly speaking, the crust-core transition takes place at a (baryon) density of $n\approx 0.096$ $\text{fm}^{-3}$ [286, 249, 127], i.e. $\rho\approx 10^{14}$ g/cm3. Crust-ocean and ocean-atmosphere transitions occur at much lower densities which are highly-temperature and composition dependent; [see Figure 2 in 179, for instance].

### v3

@@ II.1.3 SLy family
The SLy EOS is based on a semi-empirical approach that parameterises the interactions between nucleons (protons and neutrons) in terms of density-dependent coefficients of the nuclear interactions [78, 79, 128, 178]. It is derived from the Lyon-Skyrme energy-density functional (SLy) and uses a parameter set of coefficients (e.g. terms weighting the kinetic energy of nucleons) that are chosen to describe the properties of both symmetric nuclear and neutron-rich matter through fittings to six nuclear masses (spherical nuclei with an even number of neutrons and even number of protons) accounting for various properties including binding energies, charge radii, and neutron-skin thicknesses. Neutron star models constructed using the SLy EOS can reach a maximum mass of $\sim 2.1M_{\odot}$ and have a radius of $11\lesssim R\lesssim 13$ km for typical stars with 1.4 solar masses; it is therefore a relatively stiff EOS. The SLy EOS is designed to be valid for a wide range of densities, ranging from the crust up to a few times the nuclear saturation density.
Although the SLy EOS can accommodate existing mass and radius constraints and is widely used in simulations of supernovae and neutron star mergers, especially updated versions with finite-temperature corrections [178, 363, 339], it does not allow for the direct Urca or other “enhanced cooling” processes to operate. This places it in strong tension with observations of thermal emissions from accreting and isolated neutron stars [268].
@@ II.1.4 BSk family
The BSk EOS is developed through the use of the Brussels-Skyrme (BSk) energy-density functionals [^3] The BSk EOS table can be generated by the resources provided in https://www.ioffe.ru/astro/NSG/BSk/ or the CompOSE catalogue.. Like SLy, BSk is based on Skyrme-type functionals that use density-dependent coefficients to parameterise nucleon-nucleon interactions. These functionals are fitted to match experimental data on finite nuclei and nuclear matter, resulting in a sequence of incrementally named models [321, 319, 331, 171, 320]. The BSk functionals are, however, fitted to a much larger set of experimental nuclear data relative to that of SLy, namely all nuclear masses with $Z,N\geq 8$ with a model root-mean square deviation of $\approx 0.5$ MeV. The deviation from mass data (considering the restricted set of even-even nuclei) for the SLy EOS is an order of magnitude larger [$\approx 5.1$ MeV; 124]. In general, fitting terms using the largest available sets of nuclei, even and odd, spherical and deformed, is crucial to properly calibrate the functional(s).
In contrast to some other EOS, members of the family need not assume strict beta equilibrium; see Fantina et al. [138] for a description of accreted crusts with nuclear ashes for BSk19–21. Finite-temperature corrections are included but only for the surface layers at densities below $10^{6}\text{ g cm}^{-3}$. The predicted mass-radius relationship varies depending on the model number, but typically one has a radius of 11–13 km for 1.4 $M_{\odot}$ neutron stars, as for SLy. The BSk models also predict a relatively high maximum mass of between 2 and 2.4 $M_{\odot}$ In Fig. 2 we show the $M$-$R$ relation for one member of this family, BSk25, for which the maximum mass is 2.23 $M_{\odot}$ and the radius of the model with $1.4\,M_{\odot}$ is $\approx$ 12.3 km. BSk21-25 — unlike SLy4 as well as BSk19,20,26 — does allow for the direct Urca process in dense matter and conforms better to cooling observations [268].
@@ II.2 Rotation and binary alignment
Angular momentum is ubiquitous in nature. In the case of neutron stars, approximate angular momentum conservation during the supernova process implies that proto-stars can be born spinning fairly quickly, even for relatively slow progenitors. Depending on the progenitor and the impact of fallback accretion [382], some neutron stars may be born with millisecond spin periods [143]; though see also Varma et al. 428, Varma and Müller 427.
@@
Even for “slow” objects (e.g. magnetars performing a full revolution at rates less than once per few seconds), it is essential to account for rotation when considering electromagnetic observables. Depending on the rotational phase the signal can be broadened or lost due to beaming effects, which impacts interpretation. Observationally, neutron stars seem to be anything between practically static up to spins of $\sim 800$ Hz. Why there is such an observational upper limit is a topic of active research, invoking various explanations from GW-enhanced spindown to centrifugal propellering either from a companion or fallback [see, e.g. 317, for a discussion].
@@
One would like to know what the likelihood of having a “rapidly” rotating star taking part in a merger is. A review of spin properties in binaries, and the formation of double neutron-star systems in general, can be found in Tauris et al. 409. The situation is somewhat complicated however, and it has been argued that there may be three sub-populations of double neutron-star systems with distinct spin distributions, pertaining to (i) short-period and low-eccentricity binaries, (ii) wide binaries, and lastly (iii) short-period, high-eccentricity binaries [36]. The canonical formation channel involves a symbiotic binary (i.e. without dynamical captures) where there are two supernovae. These supernovae are separated by a timescale that is sensitive to a number of evolutionary specifics related to common-envelope evolution and how susceptible the companion is to Roche-lobe overflow, which in turn depends primarily on the mass ratio [e.g. 426].
@@
Prior to double neutron-star formation, accretion by the first born from the non-degenerate star can be either of a disk or wind-fed nature, and will tend to spin-up the neutron star thereby leading to a “recycled” object with a spin that is (close to) aligned with the orbital angular momentum. A discussion on predictions for alignment angles can be found in Section 2.3 of Kuan et al. 241 and references therein: several binaries have non-negligible misalignment angles as measured through geodetic precession and optical polarimetric measurements [e.g. PSR B1534+12 in a double neutron-star system with a misalignment of $\sim 27^{\circ}$; 147]. Such misalignment is relevant for the excitation of non-axisymmetric modes during inspiral (see Sec. IV.3). Either way, for a spin-up rate of $\dot{\nu}\gtrsim 10^{-14}\text{ Hz s}^{-1}$, not unusual from observations or torque theory [167], the first-born neutron star could attain spins of $\gtrsim 100$ Hz within significantly sub-Hubble times $\lesssim 10^{2}$ Myr [see table 1 in 409]. Given that the dipolar magnetic field may be buried by large factors via epochs of mass infall [278, 432, 402, 403, see also Sec. II.3], the spindown that sets in, once the secondary collapses and stops gifting material, may not be sufficient to erase spinup before merger.
@@
Zhu et al. 464 estimate that between $15$ and $30\%$ binary neutron-stars will have spins measurable via GWs at $\gtrsim 90\%$ confidence. This implies the importance of accounting for spin in tidal modelling, as all mode eigenfrequencies are skewed by even a modest degree of rotation; see equation (24) and also Appendix A of Suvorov et al. 400.
@@ II.3 Magnetic fields
Although a complete survey of how the magnetic field impresses on observables associated with the neutron star population at large lies well beyond the scope of this Review (see Enoto et al. [134] for a thorough exposition), many observations have proven that magnetic multipoles and topologically complicated structures are a reality:
@@
    The morphology of pulsed emissions are highly varied, with some systems displaying long term epochs of nulling or interpulses. Interpulse phenomenology in radio pulsars can be qualitatively explained by an oblique rotator with a multipolar magnetic field, as the emissions are then composed of multiple components [46]. The X-ray light curve from the magnetar SGR1900+14 also displays interpulse-like phenomena, which may be contributed by multiple hotspots on the surface [52]. Zhang et al. [462] suggest that starspots (i.e. localised, multipolar fields) may emerge through Hall evolution near the poles of neutron stars that are hovering around the death line, sporadically allowing the hosts to pulse and possibly explaining nulling [see also 401].
  • $\boldsymbol{\star}$
    Multipoles are typically generated via cascade phenomena through Hall drift [401] or plastic flows [173] in neutron star crusts if the “magnetic Reynolds number” is sufficiently large [$B\gtrsim\text{few}\times 10^{12}$ G; 174], and likely also through ambipolar diffusion in the core [180]. For crustal fields with a high-degree of relative helicity inverse cascades can instead operate [109], where energy is transferred from small to large scales, though this presupposes an initially multipolar field. Either way, Hall-Ohm simulation outputs are supported by observations of neutron-star cooling and (more speculatively) field decay.
@@
    Current bundles injected into the magnetosphere by crustal motions twist the fields there, inducing multipolarity [49]. Models of neutron star activity often invoke crustal failures as seeding events for outbursts [such as glitches or flares; 48, 395, 102], and therefore such injections may be common. In fact, crust failures are one of the mechanisms proposed for premerger precursors; see Sec. VI.
Aside from poloidal multipoles, the neutron star magnetic field also likely contains a multipolar and geometrically-complicated toroidal component:
@@
    Precession in magnetars, such as 4U 0142+61 [264] and SGR 1900+14 [265], are most straightforwardly explained through a (sub-)crustal toroidal field of strength $\gtrsim 10^{16}$ G: such a field introduces a prolate distortion along the magnetic axis, which then becomes misaligned with the rotation axis, causing free precession [see also 270, 396].
@@
    Mixed fields are necessary for the stability of the star. Magnetohydrodynamic (MHD) models and stability analyses demonstrate that purely poloidal/toroidal fields are unstable over dynamical timescales [411, 14, 188].
@@
    Dynamo models aiming to explain the intense fields of magnetars rely on poloidal-toroidal amplification cycles. More generally, differential rotation will wind up poloidal lines. Given that some differential rotation is inescapable as the spindown torque applied to the surface cannot instantaneously influence the core [i.e. there is a crust-core lag; 277], toroidal components should be generated.
  • $\boldsymbol{\star}$
    A universal feature of magnetothermal evolutions is that energy swaps between the toroidal and poloidal sectors, at least for sufficiently strong fields [e.g. 329, 111]. Multipolarity is thus inevitably generated in the toroidal sector from each of the considerations listed above.
@@
Figure 3 illustrates the conventional picture of pulsar operation and the presence of magnetic substructure generally (including the likelihood of internal multipoles and a superconductor; see Sec. III.4). Interpretation of neutron-star phenomena is complicated by the fact that emissions must traverse the interstellar (or intergalactic) medium, which can lens photons or GWs [see, e.g. 390].
@@
If we ignore plastic flows [though see, e.g. 174, 173, 405], the overall crustal field strength evolution may be written, after solving a simplified volume-averaged version of the induction equation following Aguilera et al. [12], as
@@
with $\tau_{\rm H}=4\pi en_{e}L^{2}/cB_{0}$ and $\tau_{\Omega}=4\pi\sigma_{\rm e}L^{2}/c^{2}$ for length-scale $L\lesssim R$ (smaller for higher-order multipoles), initial field strength $B(0)=B_{0}$, electron number-density $n_{e}$, and electrical conductivity parallel to the magnetic field $\sigma_{\rm e}$. For typical parameter choices in the crust, we may have therefore that $\tau_{H}\sim\text{few}\times 10^{4}(B_{0}/10^{14}\text{ G})^{-1}$ yr and $\tau_{\Omega}\sim\text{few}\times$ Myr. Given, however, that the GW inspiral time is orders of magnitude longer than either of these times unless the field strength is $\sim 10^{10}$ G, one anticipates weak fields at time of merger. This has important implications for the precursor flares we discuss in Sec. V.
It is important to note however that there are some ways around the decay issue, such as via Hall-stalling [i.e. the system reaches a quasi-equilibrium where $\tau_{\rm H}$ tends to some much larger number; 172], an absence of crustal impurities [which increase the conductivity and hence the Ohmic time; 203], plastic flows [which can work to halt electron flows and thus suppress cascading and the formation of multipoles; 173, 405], or dynamical capture (thereby avoiding issues related to decay altogether). Another interesting possibility relates to superconductivity (see Sec. III.4). A strong core field may be retained if the decay timescale lengthens as the star cools [196], while a strong crustal field could exist if the expulsion time of core flux into the crust is long. Whether it is at all possible for magnetar-level fields to persist over cosmological timescales remains an active area of research [see, e.g. 203, 392, for discussions].
@@ III.1 Basic elements of crustal physics
Although a neutron star may be born with temperatures in excess of $\gtrsim$ MeV, especially in a merger, the system begins to cool rapidly through a sequence of beta decays (Urca reactions) as neutrinos flood outward from what is currently an envelope [e.g. 133, 353, 365]. The details of these very early stages in the star’s life are a matter of active research and lie beyond the scope of this Review; see, for instance, Sarin and Lasky [359]. Still, this early behaviour is important as concerns some early post-merger phenomena within our purview, so some details are touched on in subsequent sections.
@@
where $Z$ denotes the number of protons in the ion, $a_{i}$ is the ion sphere radius (i.e. the Wigner-Seitz cell radius) such that $34\pi a_{i}^{3}/$ 3 equals the volume per ion, $1/n_{i}$, $e$ is the elementary charge, $k_{B}$ is the Boltzmann constant, and finally $T$ represents the temperature. Aside from fundamental constants, each ingredient defining $\Gamma$ varies with depth in a complicated way [81]. Once the temperature drops sufficiently ($\sim 10^{8}$ K) such that ${\Gamma}\gtrsim 175$ [e.g. 140], the liquid envelope solidifies via a first-order phase transition into an elastic material — the crust.
In general, because of the density dependence in expression (2), the crust may not entirely encompass the final $\gtrsim 1$ km of the star. It is expected at least that there will be an ‘ocean’ separating the crust from the magnetosphere [e.g. 164, see also Sec. VI.4.3], the physics of which depends strongly on temperature, meaning first of all that a zero-temperature EOS cannot apply [176] but also that no stress can be supported there. The crust-core transition takes place at a (baryon) density, $n$, that depends on the EOS [298, 258, 128]; for instance, $0.07\lesssim n\lesssim 0.09$ $\text{fm}^{-3}$ for the BSk EOS [320], i.e. $\rho\sim 10^{14}$ g/cm3 within a factor of a few. Crust-ocean and ocean-atmosphere transitions occur at much lower densities which are highly temperature and composition dependent [see Figure 2 in 185, for instance].